Part A : A storage tank 28 m high is filled with gasoline. (Assume the tank is open and exposed to the atmosphere at the top.)

(a) Find the gauge pressure at the bottom of the tank.( answer in Pa )

(b) Calculate the magnitude of the net force that acts on a square access hatch at the bottom of the tank that measures 0.6 m by 0.6 m.(answer in N )

Part B : What substance is it, most likely?

lead

juniper wood    

diamond

silver

brass
Answer #1

Given that

Height of storage tank h = 28 m

Using Gauge pressure formula

P =P0 + Dgh

Where

P0 = atmospheric pressure = 1.013*10^5

D is the density of gasoline D = 719.7 kg/m^3

P = 1.013*10^5 + 719.7*9.8*28 = 2.98*10^5 N /m^2

P = 2.98*10^5 pa

Part(2)

Magnitude of net force

F = pA

A is the area of the bottom tank

F =2.98*10^5*(0.6)^2

F =107280 N = 1.07*10^4 N

Part(B)

It should be silver